Alex is 4 years older than mark. Mark's age is 4/5 Alex's age. What are their ages?
Answer by stanbon(75887)

Alex is 4 years older than mark. Mark's age is 4/5 Alex's age. What are their ages?
------
Equations:
a = m + 4
m = (4/5)a
----
Substitute for "m" and solve for "a":
a = (4/5)a + 4
(1/5)a = 4
a = 20 yrs (Alex's age)
----
Solve for "m":
m = (4/5)a = (4/5)20 = 16 yrs (Mark's age)
=====================
Cheers,
Stan H.